# Towards AI-Driven Policing: Interdisciplinary Knowledge Discovery from Police Body-Worn Camera Footage

### v1

@@ Abstract
This paper proposes a novel interdisciplinary framework for analyzing police body-worn camera (BWC) footage from the Rochester Police Department (RPD) using advanced artificial intelligence (AI) and statistical machine learning (ML) techniques. Our goal is to detect, classify, and analyze patterns of interaction between police officers and civilians to identify key behavioral dynamics, such as respect, disrespect, escalation, and de-escalation. We apply multimodal data analysis by integrating video, audio, and natural language processing (NLP) techniques to extract meaningful insights from BWC footage. We present our methodology, computational techniques, and findings, outlining a practical approach for law enforcement while advancing the frontiers of knowledge discovery from police BWC data.
@@ I Introduction
Police body-worn cameras (BWCs) are expanding video-gathering tools used for evidence collection that allow researchers, crime analysts, and academics to extract insights into police dynamics in real-time. The extensive volume of data generated by BWC footage presents significant challenges across the criminal justice system, complicating efforts to analyze this information effectively and assess police-civilian interactions [1]. To address these challenges, this paper introduces the OpenBWC (Open Body-Worn Camera) framework, an open-source, multimodal system designed to leverage artificial intelligence (AI) for analyzing patterns of behavior captured in BWC footage.
     Using open-source AI technologies in a responsible and unbiased manner, this framework aims to support informed decision-making in policing practices. Our approach integrates advanced speaker separation, audio transcription, and natural language processing (NLP) techniques to examine police-civilian interactions, enabling systematic analysis and classification of behavioral patterns. By employing these methods, we aim to first classify police-civilian encounters based on key elements such as de-escalation, professionalism, and conflict resolution. Establishing this classification framework will allow us to identify critical factors influencing these outcomes. The goal is to translate analytical findings into practical strategies to improve police performance, with potential training approaches informed by collaboration with the Rochester Police Department (RPD). Current efforts focus on developing reliable methods to classify police-civilian encounters based on key behavioral and contextual elements. Consequently, this study investigates how multimodal AI techniques can systematically analyze and classify behavioral patterns in BWC footage, providing a structured foundation to improve police training, police-civilian interactions, and everyday policing practices. This approach also considers the impact of civilian behavior or error, aiming to advance public safety, equity, and the broader pursuit of justice.
@@ II Related Work
Researchers have used police BWC footage as data to learn about policing for several years already (Makin et al., 2018) [2]. With advancements in AI and NLP, this research has evolved to support large-scale analysis of recorded interactions. Computational methods have examined officer behavior, communication strategies, and officer-citizen dynamics, offering valuable contributions to policing research.
     Camp and Voigt (2025) [3] proposed a framework that applies NLP techniques to BWC footage to identify behavioral patterns, assess communication styles, and evaluate policing practices. Their approach emphasizes scaling up analysis to highlight conversational themes, officer respect, and situational factors contributing to escalation or de-escalation. While they describe how supervised and unsupervised NLP models could extract linguistic features such as politeness, apology, and reassurance, their work remains largely theoretical, with the framework’s practical implementation and performance yet to be demonstrated. While their framework shows a potential for automating behavioral analysis, the authors note that its accuracy may depend on factors such as the quality of transcriptions, the reliability of the model, and the complexity of conversations.
Additionally, they point out that noisy environments and different ways of speaking may create challenges, highlighting the importance of thorough testing in real-world policing situations. This framework’s ability to automate analysis has the potential to illuminate the nature of interactions that are often challenging to observe manually.
@@
     Moreover, recent efforts to improve the analysis of BWC footage have led to the development of video-based systematic social observation (VBSSO) methods. McCluskey et al. (2023) introduced VBSSO as a structured approach for studying police-civilian interactions using video data from police archives [6]. Their method includes strategies for sampling relevant BWC files by linking them to a department’s record management system (RMS), allowing for more efficient data selection. These techniques have influenced our framework, improving its ability to analyze behavioral patterns in police-civilian interactions.
@@ III Methodology
     At the core of the OpenBWC system is the knowledge extraction algorithm (Algorithm 1), which provides a step-by-step procedure for processing complex BWC footage. By leveraging advanced audio processing, transcription models, and NLP techniques, this algorithm extracts meaningful observations at each stage. In parallel, knowledge extraction runs alongside every pipeline component to enrich the data with semantic labels and additional insights.
Data. The dataset used in this research comprises 834 FOIL (Freedom of Information Law) videos obtained from publicly available sources through FOIL requests. These recordings capture a broad spectrum of interaction contexts, including police-civilian conversations, emergency responses, and high-pressure situations. Each BWC recording, stored as MP4 files, was systematically extracted from storage systems for further processing, including audio enhancement, transcription, and structured data organization to enable systematic and thorough analysis.
@@
Year No. of Videos Avg. Length (min)
2017 21 50.95
2019 43 5.78
2022 219 16.72
2023 365 13.62
2024 186 41.34
Total 834 128.41
Table I summarizes the distribution of videos by year, presenting the number of recordings and their average lengths. The data exhibit substantial variation, with 2024 recordings showing the longest average durations and mid-period years reflecting shorter interactions. This spread reflects a wide range of police-civilian interactions over time, offering a strong basis for analyzing contextual trends in BWC footage.
@@ IV Evaluation
The selected tools—particularly WhisperAI and SepReformer—were tested on challenging BWC footage containing background noise, overlapping speech, and varied speaker dynamics. To assess transcription quality, we manually reviewed a subset of BWC footage from routine traffic stops and compared the output transcripts with the separated audio streams. This review focused on evaluating both the accuracy of individual phrases and the clarity of speaker attribution. We found that the majority of the transcribed content—particularly officer speech—was reliably captured. This aligns with expectations, as the officer’s proximity to the BWC microphone often results in clearer audio input.
     Nonetheless, we noted instances of minor issues, including repeated phrases, occasional misinterpretations, and contextually inconsistent segments—particularly during moments of overlapping speech or increased background noise. Despite these limitations, the tools demonstrated strong baseline performance in structured interactions such as routine traffic stops.
     To further support interpretability, we applied the Llama 3.3 language model to summarize cleaned transcripts of police-civilian interactions. We guided the model to append relevant tags at the end of each summary, specifically indicating the presence of use of force, escalation, or violence. This helped us quickly identify key themes while keeping the context of each event. The model also merged the separate transcripts into one clear, speaker-labeled version, which made it easier to follow what happened. While these outputs are not a replacement for human judgment, they offer a practical starting point for downstream analysis, review, and tagging.
@@ V Limitations
    Noise Distortion. BWC footage often contains significant background noise, including sirens, vehicle sounds, and overlapping speech. Although noise reduction methods made audio clearer, sometimes too much filtering removed important details like how a person sounded or subtle cues in their voice. While text transcriptions capture the words spoken, vocal details like tone and intonation carry important meaning, emotions, and intentions—details that are often lost with too much audio filtering. Finding the right balance between reducing noise and keeping these vocal details is essential for accurately understanding noisy police encounters. Additionally, the officer’s voice is often captured with greater clarity than that of civilians, likely due to the proximity of the BWC’s microphone and its orientation toward the officer’s speech. This asymmetry in audio capture can reduce the transcription quality for civilian speech in dynamic or noisy settings.
@@
    Speaker Overlap. Overlapping speech in BWC recordings complicated transcription accuracy. Although the SepReformer model effectively separated speaker audio when only two speakers were present, it struggled in chaotic scenarios characterized by rapid interruptions or when three or more speakers overlapped. This issue impacts transcription quality during precisely the types of incidents that police managers prioritize for review—high-stakes, complex interactions. This finding aligns with previous research, such as Willits and Makin (2017) [12], who reported that most of the use-of-force cases recorded by BWCs lacked sufficient audio and video clarity for human coders to reliably interpret events. Improving audio separation techniques for multi-speaker situations represents an important area for future work, given their significance in understanding police-citizen encounters.
@@ VI Future Work
Future work will expand the multi-modal approach to include voice stress detection, sentiment analysis, and object detection to identify emotionally charged segments for more targeted manual review in high-stakes scenarios. Improved diarization models may also enhance speaker identification accuracy in complex interactions. We also plan to develop cross-referencing capabilities that allow for linking multiple incidents, enhancing pattern detection, and enabling comparison across recordings of the same event captured from different BWCs. This will help address challenges posed by concatenated video releases and support a more accurate, multi-perspective understanding of dynamic police-civilian encounters. Collaboration with police officers will further support the interpretation of algorithm-clustered incidents by grounding findings in field expertise.
@@ VII Conclusion
We introduced a framework that demonstrates the potential of combining AI-driven audio, text, and image processing with advanced AI techniques for knowledge discovery from BWC footage. Analyzing repositories of BWC footage does not require access to proprietary software; instead, the tools can be open source, and the resulting algorithms, built to analyze publicly owned data, can also remain public. This promotes transparency, reproducibility, and community engagement in future research. Building on the current analysis of a subset of approximately 834 videos, we identified key gaps—particularly in reliably assigning speaker identity during overlapping speech and maintaining transcription accuracy in high-noise, high-interruption scenarios. These limitations introduce uncertainty in the transcribed content. Addressing these gaps will be central to improving confidence in the outputs.
@@ Acknowledgment
This research is supported by a grant from the U.S. Department of Justice (DOJ) through its Office of Justice Programs (OJP) and Bureau of Justice Assistance (BJA), awarded to the Rochester Police Department (RPD).

### v2

@@ Abstract
This paper proposes a novel interdisciplinary framework for analyzing police body-worn camera (BWC) footage from the Rochester Police Department (RPD) using advanced artificial intelligence (AI) and statistical machine learning (ML) techniques. Our goal is to detect, classify, and analyze patterns of interaction between police officers and civilians to identify key behavioral dynamics, such as respect, disrespect, escalation, and de-escalation. We apply multimodal data analysis by integrating image, audio, and natural language processing (NLP) techniques to extract meaningful insights from BWC footage. We present our methodology, computational techniques, and findings, outlining a practical approach for law enforcement while advancing the frontiers of knowledge discovery from police BWC data.
@@ I Introduction
Police body-worn cameras (BWCs) are expanding video-gathering tools used for evidence collection that allow researchers, crime analysts, and academics to extract knowledge about police dynamics in real-time. The extensive volume of data generated by BWC footage presents significant challenges across the criminal justice system, complicating efforts to analyze this information effectively and assess police-civilian interactions [1]. To address these challenges, this paper introduces the OpenBWC (Open Body-Worn Camera) framework, an open-source, multimodal system designed to leverage artificial intelligence (AI) for analyzing patterns of behavior captured in BWC footage.
     Using open-source AI technologies in a responsible and unbiased manner, this framework aims to support informed decision-making in policing practices. Our approach integrates advanced speaker separation, audio transcription, and natural language processing (NLP) techniques to examine police-civilian interactions, enabling systematic analysis and classification of behavioral patterns. By employing these methods, we aim to first classify police-civilian interactions based on key elements such as de-escalation, professionalism, and conflict resolution. Establishing this classification framework will allow us to identify critical factors influencing these outcomes. The goal is to translate analytical findings into practical strategies to improve police performance, with potential training approaches informed by collaboration with the Rochester Police Department (RPD). Current efforts focus on developing reliable methods to classify police-civilian interactions based on key behavioral and contextual elements. Consequently, this study investigates how multimodal AI techniques can systematically analyze and classify behavioral patterns in BWC footage, providing a structured foundation to improve police training, police-civilian interactions, and everyday policing practices. This approach also considers the impact of civilian behavior or error, aiming to advance public safety, equity, and the broader pursuit of justice.
@@ II Related Work
Researchers have used police BWC footage as data to learn about policing for several years already (Makin et al., 2018) [2]. With advancements in AI and NLP, this research has evolved to support large-scale analysis of recorded interactions. Computational methods have examined officer behavior, communication strategies, and officer-civilian dynamics, offering valuable contributions to policing research.
     Camp and Voigt (2025) [3] proposed a framework that applies NLP techniques to BWC footage to identify behavioral patterns, assess communication styles, and evaluate policing practices. Their approach emphasizes scaling up analysis to highlight conversational themes, officer respect, and situational factors contributing to escalation or de-escalation. While they describe how supervised and unsupervised NLP models could extract linguistic features such as politeness, apology, and reassurance, their work remains largely theoretical, with the framework’s practical implementation and performance yet to be demonstrated. While their framework shows a potential for automating behavioral analysis, the authors note that its accuracy may depend on factors such as the quality of transcriptions, the reliability of the model, and the complexity of conversations. Additionally, they point out that noisy environments and different ways of speaking may create challenges, highlighting the importance of thorough testing in real-world policing situations. This framework’s ability to automate analysis has the potential to illuminate the nature of interactions that are often challenging to observe manually.
@@
     Moreover, recent efforts to improve the analysis of BWC footage have led to the development of video-based systematic social observation (VBSSO) methods. McCluskey et al. (2023) [6] introduced VBSSO as a structured approach for studying police-civilian interactions using video data from police archives. Their method includes strategies for sampling relevant BWC files by linking them to a department’s record management system (RMS), allowing for more efficient data selection. These techniques have influenced our framework, improving its ability to analyze behavioral patterns in police-civilian interactions.
@@ III Methodology
     At the core of the OpenBWC system is the knowledge extraction algorithm (Algorithm 1),
@@
Year Total Videos Lowest Range Highest Range Average Time Total Time
2021 228 00:00:30 11:58:30 3:53:36 887:39:18
2022 315 00:00:11 09:17:29 0:56:28 296:25:30
2023 352 00:00:15 11:59:30 0:59:45 350:33:38
2024 302 00:00:47 11:53:49 0:52:14 262:56:11
2025 28 00:01:34 11:29:10 2:51:59 80:15:24
Total 1225 00:00:11 11:59:30 1:31:59 1877:50:01
which provides a step-by-step procedure for processing complex BWC footage. By leveraging advanced audio processing, transcription models, and NLP techniques, this algorithm extracts meaningful observations at each stage. In parallel, knowledge extraction runs alongside every pipeline component to enrich the data with semantic labels and additional contextual information.
Data. The dataset used in this research comprises 1,225 FOIL (Freedom of Information Law) videos obtained from publicly available sources through FOIL requests. These recordings capture a broad spectrum of interaction contexts, including police-civilian conversations, emergency responses, and high-pressure situations. Each BWC recording, stored as MP4 files, was systematically extracted from storage systems for further processing, including audio enhancement, transcription, and structured data organization to enable systematic and thorough analysis.
     Table I summarizes the distribution of videos by year, presenting key duration statistics. Notably, the dataset spans recordings as short as 11 seconds to as long as nearly 12 hours. The average duration per video varies significantly by year—ranging from approximately 52 minutes in 2024 to nearly 4 hours in 2021—indicating differing operational contexts and recording practices over time. While 2021 holds the longest average video length, 2024 includes a larger number of shorter encounters, contributing to a relatively high total duration. In contrast, 2022 and 2023 exhibit more standardized recording lengths, while 2025’s smaller video count includes longer individual segments on average.
     Overall, the dataset contains a combined total duration of over 1,877 hours of footage, offering a robust foundation for identifying trends and extracting behavioral and contextual signals from a wide range of police-civilian interactions.
@@ IV Evaluation
IV-A Transcription Quality Assessment
The selected tools—particularly WhisperAI—were tested on challenging BWC footage containing background noise, overlapping speech, and varied speaker dynamics. To assess transcription quality, we manually reviewed a subset of BWC footage from routine traffic stops and compared the output transcripts with the separated audio streams. This review focused on evaluating both the accuracy of individual phrases and the clarity of speaker attribution. We found that the majority of the transcribed content—particularly officer speech—was reliably captured. This aligns with expectations, as the officer’s proximity to the BWC microphone often results in clearer audio input.
Nonetheless, we noted instances of minor issues, including repeated phrases and occasional misinterpretations—particularly during moments of overlapping speech or increased background noise.
IV-B Speaker Attribution and Role Labeling
Speaker attribution was supported by SepReformer, which helped separate overlapping voices in the BWC audio. We assessed the alignment of the separated streams with actual speakers through side-by-side comparison. The officer’s speech was generally more distinguishable and accurately attributed, again due to microphone proximity. Civilian speech occasionally required post-processing checks, particularly when multiple speakers were present or talking over one another.
IV-C Summary Tagging with LLMs
To support interpretability and thematic review, we applied the Llama 3.3 language model to summarize the cleaned transcripts of police-civilian interactions. The model was guided to append relevant tags at the end of each summary, specifically indicating the presence of use of force, escalation, or violence. This helped us quickly identify key themes while preserving the context of each event.
The model also merged the separate transcripts into one clear, speaker-labeled version, which made it easier to follow what happened. While these outputs are not a replacement for human judgment, they offer a practical starting point for downstream analysis, review, and tagging.
@@ V Limitations
    Noise Distortion. BWC footage often contains significant background noise, including sirens, vehicle sounds, and overlapping speech. Although noise reduction methods made audio clearer, sometimes too much filtering removed important details like how a person sounded or subtle cues in their voice. While text transcriptions capture the words spoken, vocal details like tone and intonation carry important meaning, emotions, and intentions—details that are often lost with too much audio filtering. Finding the right balance between reducing noise and keeping these vocal details is essential for accurately understanding noisy police-civilian interactions. Additionally, the officer’s voice is often captured with greater clarity than that of civilians, likely due to the proximity of the BWC’s microphone and its orientation toward the officer’s speech. This asymmetry in audio capture can reduce the transcription quality for civilian speech in dynamic or noisy settings.
@@
    Speaker Overlap. Overlapping speech in BWC recordings complicated transcription accuracy. Although the SepReformer model effectively separated speaker audio when only two speakers were present, it struggled in chaotic scenarios characterized by rapid interruptions or when three or more speakers overlapped. This issue impacts transcription quality during precisely the types of incidents that police managers prioritize for review—high-stakes, complex interactions. This finding aligns with previous research, such as Willits and Makin (2017) [12], who reported that most of the use-of-force cases recorded by BWCs lacked sufficient audio and video clarity for human coders to reliably interpret events. Improving audio separation techniques for multi-speaker situations represents an important area for future work, given their significance in understanding police-civilian interactions.
@@ VI Future Work
Future work will expand the multi-modal approach to include voice stress detection, sentiment analysis, and object detection to identify emotionally charged segments for more targeted manual review in high-stakes scenarios. Improved diarization models may also enhance speaker identification accuracy in complex interactions. We also plan to develop cross-referencing capabilities that allow for linking multiple incidents, enhancing pattern detection, and enabling comparison across recordings of the same event captured from different BWCs. This will help address challenges posed by concatenated video releases and support a more accurate, multi-perspective understanding of dynamic police-civilian interactions. Collaboration with police officers will further support the interpretation of algorithm-clustered incidents by grounding findings in field expertise.
@@ VII Conclusion
We introduced a framework that demonstrates the potential of combining AI-driven audio, text, and image processing with advanced AI techniques for knowledge discovery from BWC footage. Analyzing repositories of BWC footage does not require access to proprietary software; instead, the tools can be open source, and the resulting algorithms, built to analyze publicly owned data, can also remain public. This promotes transparency, reproducibility, and community engagement in future research. Building on the current analysis of 1,225 videos, we identified key gaps—particularly in reliably assigning speaker identity during overlapping speech and maintaining transcription accuracy in high-noise, high-interruption scenarios. These limitations introduce uncertainty in the transcribed content. Addressing these gaps will be central to improving confidence in the outputs.
@@ Acknowledgment
This research is supported by a grant from the U.S. Department of Justice (DOJ) through its Office of Justice Programs (OJP) and Bureau of Justice Assistance (BJA), awarded to the RPD.